OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = 3x + 4. What is f(g(x)) ?

jimthompson5910 (jim_thompson5910):

f(x) = x – 2 f(g(x)) = g(x) – 2 ... replace every x with g(x) see what to do now?

OpenStudy (anonymous):

3x+2=f(x)

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

no i'm not sure how you're getting that

OpenStudy (anonymous):

You said replace x with the value of g(x)

OpenStudy (anonymous):

So I did that and simplified,

OpenStudy (anonymous):

Can you help DLS?

jimthompson5910 (jim_thompson5910):

f(g(x)) = g(x) – 2 Now replace g(x) on the right side with 3x+4 (since they are the same)

jimthompson5910 (jim_thompson5910):

then simplify

OpenStudy (anonymous):

That's what I did. 3x+2

jimthompson5910 (jim_thompson5910):

oh ok, your notation above is a bit off up there

jimthompson5910 (jim_thompson5910):

it should be f(g(x)) = 3x+2

OpenStudy (anonymous):

Ya. I forgot the g.

jimthompson5910 (jim_thompson5910):

thats ok

OpenStudy (anonymous):

:3